# Do Prompts Really Prompt? Exploring the Prompt Understanding Capability of Whisper

### v2

@@ 1 Introduction
Research in NLP [37] has shown that models can be enhanced by prompting even without fully understanding the prompts, suggesting that performance improvements do not necessarily imply true prompt comprehension because prompts with incorrect information may also improve the model like those with correct information, which is not expected for a model with true prompt understanding. Hence, because of the domain shifts mentioned above, the assumption that Whisper understands prompts requires scrutiny despite the observed performance gains in earlier studies. Additionally, a deeper understanding of the prompting mechanism provides better insights into Whisper, facilitating further development and application.
This work explores Whisper’s prompt understanding by analyzing its performances when prompted with information that matches or mismatches with the testing data in several scenarios. Ideally, if the model understands the prompts, matched prompts should lead to better performance. Our systematic analysis based on three novel metrics reveals that Whisper exhibits limited semantic understanding of textual prompts, often performing better with mismatched information. We also find that textual prompts hurt the performance in general. Further analysis surprisingly reveals that no positive correlation between prompt understanding and performance exists, showing that prompting Whisper does not work as expected. Additionally, prompts in different languages from the testing data often outperform those in the same language, showcasing a counter-intuitive effect of prompt language. In code-switched ASR, expected performance degradation occurs with nonexistent language token pairs, showing that Whisper can filter out irrelevant information within the language tokens, likely benefiting from extensive pre-training.
Overall, our contributions [^1] Resources of this work can be found at https://anonymous.4open.science/r/whisper_prompting-90B2. are: 1) Pioneering exploration of Whisper’s prompt understanding ability with three newly proposed metrics, 2) Uncovering the unexpected behaviors of Whisper when given textual prompts and empirically demonstrating that adhering strongly to topic information in prompts does not guarantee improved performances, 3) Revealing a counter-intuitive effect of prompt languages where English prompts outperform Mandarin ones, even with Mandarin testing data, and 4) Highlighting its ability to disregard misleading information in language tokens. These findings aim to stimulate further research on prompting methods and the development of Whisper and other universal ASR models with strong prompt understanding.
@@ 2.2 Prompting Whisper
Prompting Whisper is widely applied. It has been known that prompting can make Whisper follow a certain style [38] when transcribing, e.g., transcribing in Traditional Chinese instead of Simplified Chinese when given a prompt in Traditional Chinese, etc. However, the application of prompting Whisper is much more than this. Peng et al. [33] showed that Whisper can be prompted for audio-visual ASR by including visual objects retrieved by CLIP [39] in textual prompts. They also showed that concatenating two language tokens enhanced Whisper on code-switched (CS) ASR for those languages. Yang et al. [34] observed improvement in CS ASR by indicating code-switching occurrences in textual prompts. Wang et al. [35] showcased the efficacy of task-informed textual prompts, e.g., “recognize dialect speech”, and in-context learning in boosting Whisper on dialect ASR. Zhuo et al. [36] generalized Whisper to lyric transcription via prompting. These studies show the potential of prompting Whisper beyond style imitation. However, whether Whisper understands the prompts remains unexplored, motivating this systematic investigation of Whisper’s prompt understanding.
@@ 3.2 Method
Mismatched prompts are designed systematically to address potential confounding factors. Specifically, mismatched prompts are generated by introducing incorrect information that does not align with the testing data. For semantic understanding of textual prompts, we construct the prompt templates (Sec. 3.2.1) and introduce semantic corruption systematically (Sec. 3.2.2). The templates follow various patterns, ensuring our results are not biased by certain prompt structures. For the effect of textual prompt languages, the corruption is introduced by using prompts translated into a language different from the data while maintaining the meaning (Sec. 3.2.3), mitigating the effect of semantic variation. Regarding the language tokens, we provide a pair of language tokens of the languages occurring in the data and corrupt the pairs by replacing one token with a non-existent language (Sec. 3.2.4). No textual prompt is provided and all the other settings are identical.
Although this work focuses on Whisper, our method can be generalized to other speech models. Different models may handle prompts and mismatches differently, but our systematic approach can evaluate their prompt understanding capabilities. This generalizability ensures our method is a valuable tool for advancing prompt understanding research in speech-processing models.
@@ 3.2.1 Prompt templates for textual prompts
To explore Whisper’s textual prompt understanding, we construct prompt templates to generate textual prompts. These templates transform inputs into strings. For instance, a template $\mathtt{This}$ $\mathtt{utterance}$ $\mathtt{is}$ $\mathtt{about}$$\{\mathtt{input}\}$ turns $\mathtt{arts}$ into a prompt “This utterance is about arts”. Using templates helps systematically generate textual prompts with the desired information.
To ensure diversity, templates are first generated manually with various patterns, where some resemble those in prior works, categorized as follows with examples: (1) identity mappings, which use the input as prompts: $\{\mathtt{input}\}$; (2) task-informed instructions [35, 36]: $\mathtt{Transcribe}$ $\mathtt{this}$ $\mathtt{video}$ $\mathtt{of}$$\{\mathtt{input}\}$; (3) conversation-like templates [38]: $\mathtt{So}$ $\mathtt{we}$ $\mathtt{were}$ $\mathtt{just}$ $\mathtt{talking}$ $\mathtt{about}$$\{\mathtt{input}\}$; (4) indication of the input information in natural language form [34]: $\mathtt{This}$ $\mathtt{utterance}$ $\mathtt{is}$ $\mathtt{about}$$\{\mathtt{input}\}$; (5) lists of input-related keywords [33] generated by GPT-3.5 [43]: $\mathtt{\{keyword\_1\},}$$\mathtt{\{keyword\_2\},...,}$$\mathtt{\{keyword\_N\}}$. In particular, human-generated templates of patterns (2), (3), and (4) are provided to GPT-3.5 to generate additional templates. All the templates are generated in English. The resulting 10 templates generated by humans and GPT-3.5 are collected, and textual prompts are subsequently constructed with the initial character capitalized. For instance, if the input is $\mathtt{arts}$, the generated prompts of the aforementioned templates are (1) “Arts”, (2) “Transcribe this video of arts”, (3) “So we were just talking about arts”, (4) “This utterance is about arts”, and (5) “Arts, culture, performing, visual”.
@@ 3.2.2 Semantic understanding of textual prompts
Since matched and mismatched prompts differ only in topic correctness, the performance gap between these scenarios indicates the model’s sensitivity to topic information and measures its semantic understanding of textual prompts. We compare performance across subsets of each topic, examining prompts with correct and mismatched topics, using metrics from Sec. 3.3. In addition, the performances under the no-prompt settings are also discussed to gain more insights.
@@ 3.2.3 Effect of textual prompt languages
We explore whether the languages of textual prompts influence the model’s performance. That is, whether Whisper gets enhanced by implicit language information in textual prompts. During the pre-training, the decoder generates predictions conditioned on the content of the previous speech segment, typically in the same languages as the testing speech segment. Consequently, textual prompts in the same language as the testing data are expected to yield better performance since it is closer to the pre-training setting. To test this hypothesis, we compare Whisper’s transcription accuracy on English and Mandarin data when prompted by English or Mandarin prompts with identical meanings.
@@ 4.2.1 Multi-domain ASR corpora
Multi-domain ASR corpora with associated topic labels are employed to investigate Whisper’s understanding of textual prompts. We incorporate GigaSpeech [44], a multi-domain English ASR corpus. We utilize the training split [^2] No models trained. Using training split for evaluation is acceptable. of its S subset, as the testing split lacks topic labels. Notably, we notice that some topics in GigaSpeech, e.g., “People and Blogs” and “Entertainment,” are quite rough, and their data may be semantically related to other topics simultaneously. To avoid potential overlap, we select subsets of four distinct topics: “Arts,” “Science and Technology,” “Nonprofits and Activism,” and “Sports,” totaling 34.4 hours.
@@ 5 Results and discussion
We discuss our results. The 95% confidence intervals obtained from bootstrapping [47], with 1k bootstrap sets, are shown in brackets. Confidence intervals for TFRs are omitted due to the limited number of templates. Each experiment takes around 10 hours on a V100.
@@ 5.1 Semantic understanding of textual prompts
From Table 1, Whisper’s average topic-following rates (TFRs) are markedly unsatisfactory on the testing corpora, typically below 30%. This indicates that in general, Whisper does not perform better when prompted by matched information.
The PERFs and BPERFs of Whisper on the evaluation corpora are analyzed. To ensure the comparison between PERFs and BPERFs can reflect the model’s semantic prompt understanding, we first compute the word error rates for non-semantic errors like function words and observe a maximum difference of only 0.5% between the predictions contributing to PERFs and BPERFs. This negligible difference suggests that such errors occur almost equally in settings with both matched and mismatched prompts, supporting our hypothesis in Section 3.3.2. Thus, the performance discrepancies between BPERFs and PERFs effectively indicate the semantic understanding of textual prompts. From Table 1, the mismatched information can induce a relative improvement of BPERF to PERF of up to 11%, which is highly significant. This means that in addition to low topic-following rates, Whisper gets greatly enhanced by irrelevant and incorrect information in textual prompts. In addition, we can see that Whisper’s average PERFs/BPERFs are typically worse than the no-prompt settings. Given that there are already 10 diverse templates resembling prior works in our experiments, which is sufficient to mitigate the potential bias of the template choices, this result indicates that textual prompts that make sense to humans do not always bring improvements.
@@ 5.2 Effect of textual prompt languages
This might be attributed to the substantial disparity in the amount of training data for English and Mandarin [21]. Whisper was primarily trained on English data, and the previous context on which the generation was conditioned was English most time. This might make Whisper more familiar with English than other languages. Thus, when replacing the previous context with textual prompts, English prompts may elicit better performance than their Mandarin counterparts, regardless of the language discrepancy between prompts and the speech.
@@ 6 Conclusion and Future Work
We study Whisper’s prompt understanding. Despite the widespread use of prompting Whisper, we discovered that informative prompts in human perception do not improve Whisper in general, and better performance is not guaranteed even when demonstrating better comprehension given a prompt template. We found that English prompts yield better performance than translated Mandarin ones on both English and Mandarin data, possibly due to differences in the training data volume for these languages. It was also noted that Whisper can filter out misleading information in language tokens, likely due to extensive pre-training on language identification and ASR. Future work includes automatically constructing effective prompts and establishing speech models with genuine prompt understanding. We encourage the community to carefully revisit Whisper’s prompting methods and provide more investigations.

### v4

@@ 1 Introduction
Research in NLP [37] has shown that models can be enhanced by prompting even without fully understanding the prompts, suggesting that performance improvements do not necessarily imply true prompt comprehension because prompts with incorrect information may also improve the model like those with correct information, which is not expected for a model with true prompt understanding. Hence, in light of the domain shifts mentioned above, the assumption that Whisper understands prompts requires scrutiny despite the observed performance gains in earlier studies. Additionally, a deeper understanding of the prompting mechanism provides better insights into Whisper, facilitating further development and application.
This work explores Whisper’s prompt understanding by analyzing its performances when prompted with information that matches or mismatches with the testing data in several scenarios. The information is encoded into prompts through special tokens or certain textual templates, and thus the “prompt understanding” is required for the model to process the information. Ideally, if the model possesses good prompt understanding, matched prompts should lead to better performance. Our systematic analysis based on three novel metrics reveals that Whisper exhibits limited semantic understanding of textual prompts, often performing better with mismatched information. We also find that textual prompts hurt the performance in general. Further analysis surprisingly reveals that no positive correlation between prompt understanding and performance exists, showing that prompting Whisper does not work as expected. Additionally, prompts in different languages from the testing data often outperform those in the same language, showcasing a counter-intuitive effect of prompt language. In code-switched ASR, expected performance degradation occurs with nonexistent language token pairs, showing that Whisper can filter out irrelevant information within the language tokens, likely benefiting from extensive pre-training.
Overall, our contributions [^1] Resources of this work can be found at https://github.com/b08202033/whisper_prompting. are: 1) Pioneering exploration of Whisper’s prompt understanding ability with three newly proposed metrics, 2) Uncovering the unexpected behaviors of Whisper when given textual prompts and empirically demonstrating that adhering strongly to topic information in prompts does not guarantee improved performances, 3) Revealing a counter-intuitive effect of prompt languages where English prompts outperform Mandarin ones, even with Mandarin testing data, and 4) Highlighting its ability to disregard misleading information in language tokens. These findings serve as solid foundations for further research and applications on prompting methods and the development of Whisper and other universal ASR models with strong prompt understanding.
@@ 2.2 Prompting Whisper
Prompting Whisper is widely applied. It has been known that prompting can make Whisper follow a certain style [38] when transcribing, e.g., transcribing in Traditional Chinese instead of Simplified Chinese when given a prompt in Traditional Chinese, etc. However, the application of prompting Whisper is much more than this. Peng et al. [33] showed that Whisper can be prompted for audio-visual ASR by including visual objects retrieved by CLIP [39] in textual prompts. They also showed that concatenating two language tokens enhanced Whisper on code-switched (CS) ASR for those languages. Yang et al. [34] observed improvement in CS ASR by indicating code-switching occurrences in textual prompts. Wang et al. [35] showcased the efficacy of task-informed textual prompts, e.g., “recognize dialect speech”, and in-context learning in boosting Whisper on dialect ASR. Zhuo et al. [36] generalized Whisper to lyric transcription via prompting. These studies show the potential of prompting Whisper beyond style imitation. However, whether Whisper understands the prompts remains unexplored, motivating our systematic investigation of Whisper’s prompt understanding.
@@ 3.2 Method
Mismatched prompts are designed systematically to address potential confounding factors. Specifically, they are generated by introducing incorrect information that does not align with the testing data. For semantic understanding of textual prompts, we construct the prompt templates (Sec. 3.2.1) and introduce semantic corruption systematically (Sec. 3.2.2). The templates follow various patterns, ensuring our results are not biased by certain prompt structures. For the effect of textual prompt languages, the corruption is introduced by using prompts translated into a language different from the data while maintaining the meaning (Sec. 3.2.3), mitigating the effect of semantic variation. Regarding the language tokens, we provide a pair of language tokens of the languages occurring in the data and corrupt the pairs by replacing one token with a non-existent language (Sec. 3.2.4). No textual prompt is provided here and all the other settings are identical.
Although this work focuses on Whisper, our method can be generalized to other speech models. Different models may handle prompts and mismatches differently, but our systematic approach can still evaluate their prompt understanding capabilities. This generalizability ensures our method is a valuable tool for advancing prompt understanding research in speech-processing models.
@@ 3.2.1 Prompt templates for textual prompts
To explore Whisper’s textual prompt understanding, we construct prompt templates to generate textual prompts. These templates transform inputs into strings. For instance, a template $\mathtt{This}$ $\mathtt{utterance}$ $\mathtt{is}$ $\mathtt{about}$$\{\mathtt{input}\}$ turns $\mathtt{arts}$ into a prompt “This utterance is about arts”. Using templates helps systematically generate textual prompts with the desired information encoded.
To ensure diversity, templates are first generated manually with various patterns, where some resemble those in prior works, categorized as follows with examples: (1) identity mappings, which use the input as prompts: $\{\mathtt{input}\}$; (2) task-informed instructions [35, 36]: $\mathtt{Transcribe}$ $\mathtt{this}$ $\mathtt{video}$ $\mathtt{of}$$\{\mathtt{input}\}$; (3) conversation-like templates [38]: $\mathtt{So}$ $\mathtt{we}$ $\mathtt{were}$ $\mathtt{just}$ $\mathtt{talking}$ $\mathtt{about}$$\{\mathtt{input}\}$; (4) indication of the input information in natural language form [34]: $\mathtt{This}$ $\mathtt{utterance}$ $\mathtt{is}$ $\mathtt{about}$$\{\mathtt{input}\}$; (5) lists of input-related keywords [33] generated by GPT-3.5 [43]: $\mathtt{\{keyword\_1\},}$$\mathtt{\{keyword\_2\},...,}$$\mathtt{\{keyword\_N\}}$. In particular, human-generated templates of patterns (2), (3), and (4) are provided to GPT-3.5 to generate additional templates. All templates are generated in English. The resulting 10 templates generated by humans and GPT-3.5 are collected, and textual prompts are subsequently constructed with initial characters capitalized. For instance, given the input: $\mathtt{arts}$, the generated prompts of the aforementioned templates are (1) “Arts”, (2) “Transcribe this video of arts”, (3) “So we were just talking about arts”, (4) “This utterance is about arts”, and (5) “Arts, culture, performing, visual”.
@@ 3.2.2 Semantic understanding of textual prompts
Since matched and mismatched prompts differ only in topic correctness, the performance gap between these scenarios indicates the model’s sensitivity to topic information and measures its semantic understanding of textual prompts. We compare performance across subsets of each topic, examining prompts with correct and mismatched topics, using metrics from Sec. 3.3. In addition, the performances under the no-prompt settings are also discussed to gain more insights into the general effectiveness of prompting.
@@ 3.2.3 Effect of textual prompt languages
We explore whether the languages of textual prompts influence Whisper’s performance, i.e., whether it gets enhanced by implicit language information in textual prompts. During the pre-training, the decoder predicts conditioned on the content of the previous speech segment, typically in the same languages as the testing segment. Consequently, textual prompts in the same language as the testing data should yield better performance since it is closer to the pre-training setting. To test this hypothesis, we compare Whisper’s transcription accuracy on English and Mandarin data when prompted by English or Mandarin prompts with identical meanings.
@@ 4.2.1 Multi-domain ASR corpora
Multi-domain ASR corpora with associated topic labels are employed to investigate Whisper’s understanding of textual prompts. We incorporate GigaSpeech [44], a multi-domain English ASR corpus. We utilize the training split [^2] No models trained. Using training split for evaluation is acceptable. Though it is unknown whether these data were used during pre-training, the boost from pre-training can be canceled when we only focus on the difference of performances in two scenarios, since such boost will affect them similarly. of its S subset, as the testing split lacks topic labels. Notably, we notice that some topics in GigaSpeech, e.g., “People and Blogs” and “Entertainment,” are quite rough, and their data may be semantically related to other topics simultaneously. To avoid potential overlap, we select subsets of four highly distinct topics: “Arts,” “Science and Technology,” “Nonprofits and Activism,” and “Sports,” totaling 34.4 hours.
@@ 5 Results and discussion
We discuss our results. The 95% confidence intervals (CIs) obtained from bootstrapping [47], with 1k bootstrap sets, are shown in brackets in the tables. CIs for TFRs are omitted due to the limited number of templates. Each experiment takes around 10 hours on a V100.
@@ 5.1 Semantic understanding of textual prompts
From Table 1, Whisper’s average topic-following rates (TFRs) are markedly unsatisfactory on the testing corpora, typically below 30%. This indicates that Whisper does not perform better when prompted by matched information in general.
The PERFs and BPERFs of Whisper on the evaluation corpora are analyzed. First of all, we notice that Whisper’s average PERFs/BPERFs are typically worse than the no-prompt settings. We further compare the PERFs and BPERFs. To ensure the comparison between PERFs and BPERFs can reflect the model’s semantic prompt understanding, we first compute the word error rates for non-semantic errors like function words and observe a maximum difference of only 0.5% between the predictions contributing to PERFs and BPERFs. This negligible difference suggests that such errors occur almost equally in settings with both matched and mismatched prompts, supporting our hypothesis in Section 3.3.2. Thus, the performance discrepancies between BPERFs and PERFs effectively indicate the semantic understanding of textual prompts. From Table 1, the mismatched information can induce relative improvements of BPERF to PERF of up to 11%, which are highly significant. Such improvements are commonly observed for templates inducing better or worse performances than the no-prompt baseline, showing that besides low TFRs, Whisper gets greatly enhanced by irrelevant and incorrect information in textual prompts. Given that there are already 10 diverse templates resembling prior works in our experiments, which is sufficient to mitigate the potential bias of the template choices, this result indicates that textual prompts that make sense to humans do not always bring improvements.
@@ 5.2 Effect of textual prompt languages
This might be attributed to the substantial disparity in the amount of training data for English and Mandarin [21]. Whisper was primarily trained on English data, and the previous context on which the generation was conditioned was typically in English. This might make Whisper more familiar with English than other languages. Thus, when replacing the previous context with textual prompts, English prompts may elicit better performance than their Mandarin counterparts, regardless of the language discrepancy between prompts and the speech.
@@ 6 Conclusion and Future Work
We study Whisper’s prompt understanding. Despite the widespread use of prompting Whisper, we discovered that informative prompts in human perception do not improve Whisper in general, and better performance is not guaranteed even when demonstrating better comprehension given a prompt template. We found that English prompts yield better performance than Mandarin counterparts on both English and Mandarin data, possibly due to differences in the training data volume for these languages. It was also noted that Whisper can filter out misleading information in language tokens, likely due to extensive pre-training on language identification and ASR. Future work includes automatically constructing effective prompts and establishing speech models with genuine prompt understanding. We encourage the community to carefully revisit Whisper’s prompting methods and provide more investigations.
7 Acknowledgement
We acknowledge the computational and storage support from National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan.